# Local Differential Privacy for Tensors in Distributed Computing Systems

### v2

@@ V-B2 Type II: Private Training Features
$s_{2}(f)=\underset{\mathcal{X},\mathcal{X}^{\prime}}{sup}||\mathcal{X}-\mathcal{X}^{\prime}||_{F}=2\sqrt{I_{1}I_{2}\dots I_{N}},$ (18)
since the range of data is set to $(-1,1)$.
TABLE III: Datasets and parameter settings for Type II
Dataset MNIST Cifar-10 SVHN
Model MobileNet MobileNet MobileNet
Training Data 60000 50000 73257
Testing Data 10000 10000 26032
Batch size 64 64 64
$\Delta$ 256 256 256
V-B3 Type III: Private SGD
For this task, we utilized datasets including MNIST, CIFAR-10, SVHN, IMDB, and DarkNet, as shown in Table IV. The training data is split into three parts, and each local model is trained on one subset. After completing a single epoch, each local model uploads its weights and biases using the global model. The global model then updates its parameters by averaging the weights and biases from the three local models and subsequently sends the updated parameters back to the local models. As a result, after each iteration, the weights and biases of all four models are synchronized. The entire process can be restated as follows:
@@ V-B3 Type III: Private SGD
$s_{2}(f)=\underset{\mathcal{X},\mathcal{X}^{\prime}}{sup}||\mathcal{X}-\mathcal{X}^{\prime}||_{F}=2C\sqrt{I_{1}I_{2}\dots I_{N}}$ (19)
@@
TABLE IV: Datasets and parameter settings for Type III
Dataset MNIST Cifar-10 SVHN IMDB DarkNet
Model MobileNet MobileNet MobileNet BiLSTM MLP
Training Data 60000 50000 73257 25000 113184
Testing Data 10000 10000 26032 25000 28297
Batch size 64 64 64 64 64
Clip value 1 1 1 1 1
$\Delta$ 256 256 256 400001 256
V-C Experimental Results
Before comparing the experimental outcomes, we provide a theoretical analysis of the expected error associated with each (local) differential privacy approach. The details of these theoretical findings are summarized in Table V.
TABLE V: COMPARISON OF EXPECTED ERROR
Method Laplace Gaussian MVG IDN TLDP-L TLDP-G
$\mathbb{E}||\mathcal{Z}||_{2}$ $\frac{\Delta}{\epsilon}I\sqrt{2I}$ $\frac{\Delta}{2\epsilon}I\sqrt{I}$ $\frac{I_{1}I}{(-\beta_{0}+\sqrt{\beta_{0}^{2}+8\alpha_{0}\epsilon})^{2}}\cdot 4\alpha_{0}^{2}$ $\sqrt{\frac{Is_{2}^{2}(f)}{(-\zeta(\delta)+\sqrt{\zeta^{2}(\delta)+2\epsilon})^{2}}}$ $\frac{\Delta}{\epsilon}\sqrt{2(1-p)I}$ $\frac{\Delta}{2\epsilon}\sqrt{(1-p)I}$
In Table V, $I=I_{1}I_{2}\dots I_{N}$ is a product of each dimension of the tensor. Since the query functions for the three tasks outlined in Section V-B are identical, the expected errors are consistent, which is defined as
$\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}||\mathcal{Z}||_{2}=\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}(\sum_{i=1}^{I}(\mathcal{Z}_{i})^{2})^{\frac{1}{2}},$ (20)
@@ V-C Experimental Results
$\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}||\mathcal{Z}||_{2}=\frac{\Delta}{\epsilon}\sqrt{2(1-p)I}.$ (22)
@@
$\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}||\mathcal{Z}||_{2}=\frac{\Delta}{2\epsilon}\sqrt{(1-p)I}.$ (23)
If the randomized response mechanism is removed from our method, i.e., no component will be retained with the probability $p$, and the term $1-p$ in the outcomes will become $1$. Moreover, as discussed in the proof of Theorem 2, in order to ensure the same privacy budget, the values of the two noise parameters are increased by a factor of $I$ compared to the improved methods. As a result, the expected errors of normal Laplace and Gaussian mechanisms are $\frac{\Delta}{\epsilon}I\sqrt{2I}$ and $\frac{\Delta}{2\epsilon}I\sqrt{I}$, respectively.
For MVG and IDN, we adopt the same parameters as Task II in [16] and [17], which provides the expected error $\frac{I_{1}I}{(-\beta_{0}+\sqrt{\beta_{0}^{2}+8\alpha_{0}\epsilon})^{2}}\cdot 4\alpha_{0}^{2}$ and $\sqrt{\frac{Is_{2}^{2}(f)}{(-\zeta(\delta)+\sqrt{\zeta^{2}(\delta)+2\epsilon})^{2}}}$ (where $\zeta^{2}(\delta)=-2\ln{\delta}+2\sqrt{-I\ln{\delta}}+I$). Removing the relaxation term $\delta$, we obtain its estimated value $\frac{\Delta I_{1}I\sqrt{I}}{\sqrt{2}\epsilon}\cdot\ln^{2}(I_{1}+1)$ and $\frac{\Delta I\sqrt{I}}{\sqrt{2}\epsilon}.$
@@ V-C2 Type II Results
Fig. 2 presents the F1-score across different values of $\epsilon$, where the relaxation term $\delta$ is set to $10^{-5}$. Compared to Type I, the rate of increase in F1-score across the three experimental results slows down as $\epsilon$ increases in Type II. For the MNIST dataset, the F1-score reaches a maximum of about 80%, while in CIFAR-10 and SVHN, the increase is much more limited, with F1-scores hovering around 40% finally. Unlike in Type I, the performance of the Laplace, Gaussian, MVG, and IDN mechanisms in Type II shows occasional fluctuations in F1-scores, although they still remain at extremely low levels, indicating them ineffective for classification tasks. Additionally, consistent with the theoretical error analysis presented earlier, the TLDP-G mechanism generally outperforms the TLDP-L mechanism in these experiments.
V-C3 Type III Results
@@
Fig. 3 shows the results of Type III on the different datasets for comparison. Two new datasets, IMDB for text mining and DarkNet for data mining, are introduced in the experiment. In Type III, due to the transmission of more abstract model parameters to the central server, the results from the various methods only become comparable when $\epsilon$ exceeds 100. As anticipated, the performance of the Laplace, Gaussian, MVG, and IDN mechanisms remains significantly low. In contrast, the TLDP method consistently delivers superior results, particularly on the MNIST, SVHN, IMDB, and DarkNet datasets.
@@
Fig. 3 displays the F-1 score for varying $\epsilon$ values, with the relaxation term $\delta$ fixed as $10^{-5}$. In Type III, the results for the MNIST, CIFAR-10, and SVHN datasets are consistent with those in Type I and Type II, with TLDP significantly outperforming the Laplace, Gaussian, MVG, and IDN mechanisms. On the IMDB dataset, the accuracy of the Laplace, Gaussian, MVG, and IDN mechanisms improves to around 40%, but no further improvement is observed as $\epsilon$ increases. In contrast, both TLDP methods consistently achieve accuracy levels of around 80%. On the DarkNet dataset, the performance of the different methods is more distinctly differentiated. The IDN and Laplace mechanisms show the poorest performance, followed by the Gaussian and MVG mechanisms. Although the TLDP method continues to demonstrate an advantage, the gap with other methods is less pronounced compared to the former experiments. This is especially true for the TLDP-G mechanism, which initially exhibits the lowest accuracy when $\epsilon$ is set to 10 but eventually surpasses all other methods, aligning with the TLDP-L curve.
We find that the performance of IDN is not as good as claimed in [17], primarily due to its high reliance on the utility subspace. When there is a lack of known conditions, its performance is significantly compromised.
@@
The results from all the aforementioned experiments provide strong evidence of the robustness of the TLDP mechanism across diverse experimental settings. Regardless of the dataset or the tasks, TLDP consistently outperforms other baseline methods, showcasing its ability to adapt effectively to different environments. This robustness highlights TLDP’s potential as a reliable solution for preserving privacy while maintaining high model performance, making it a promising candidate for a wide range of applications in real-world scenarios.
VI Conclusion
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62406215 and 62072321).

### v3

@@ 5.2.2 Type II: Private Training Features
[…]
since the range of data is set to (1,1)(-1,1).
Table 3: Datasets and parameter settings for Type II
Dataset Model Training Data Testing Data Batch size $\Delta$
MNIST MobileNet 60000 10000 64 256
CIFAR-10 MobileNet 50000 10000 64 256
SVHN MobileNet 73257 26032 64 256
5.2.3 Type III: Private SGD
For this task, we utilized datasets including MNIST, CIFAR-10, SVHN, IMDB, and DarkNet, as shown in Table 4. The training data is split into three parts, and each local model is trained on one subset. After completing a single epoch, each local model uploads its weights and biases using the global model. The global model then updates its parameters by averaging the weights and biases from the three local models and subsequently sends the updated parameters back to the local models. As a result, after each iteration, the weights and biases of all four models are synchronized. The entire process can be restated as follows:
  • 1.
@@ 5.2.3 Type III: Private SGD
[…]
@@
Table 4: Datasets and parameter settings for Type III
Dataset Model Training Data Testing Data Batch size Clip value $\Delta$
MNIST MobileNet 60000 10000 64 1 256
CIFAR-10 MobileNet 50000 10000 64 1 256
SVHN MobileNet 73257 26032 64 1 256
IMDB BiLSTM 25000 25000 64 1 400001
DarkNet MLP 113184 28297 64 1 256
5.3 Experimental Results
Before comparing the experimental outcomes, we provide a theoretical analysis of the expected error associated with each (local) differential privacy approach. The details of these theoretical findings are summarized in Table 5.
Table 5: COMPARISON OF EXPECTED ERROR
Method Laplace Gaussian MVG IDN DPHSGD TLDP-L TLDP-G
[…]
In Table 5, $I=I_{1}I_{2}\dots I_{N}$ is a product of each dimension of the tensor. Since the query functions for the three tasks outlined in Section 5.2 are identical, the expected errors are consistent, which is defined as
[…]
@@ 5.3 Experimental Results
[…]
@@
[…]
If the randomized response mechanism is removed from our method, i.e., no component will be retained with the probability $p$, and the term 1p1-p in the outcomes will become $1$. Moreover, as discussed in the proof of Theorem 2, in order to ensure the same privacy budget, the values of the two noise parameters are increased by a factor of $I$ compared to the improved methods. As a result, the expected errors of normal Laplace and Gaussian mechanisms are $\frac{\Delta}{\epsilon}I\sqrt{2I}$ and $\frac{\Delta}{2\epsilon}I\sqrt{I}$, respectively.
For MVG and IDN, we adopt the same parameters as Task II in [24] and [25], which provides the expected error
[…]
and
[…]
where $\zeta^{2}(\delta)=-2\ln{\delta}+2\sqrt{-I\ln{\delta}}+I$. Removing the relaxation term $\delta$, we obtain its estimated value $\frac{\Delta I_{1}I\sqrt{I}}{\sqrt{2}\epsilon}\cdot\ln^{2}(I_{1}+1)$ and $\frac{\Delta I\sqrt{I}}{\sqrt{2}\epsilon}$. Compared with the Gaussian mechanism, DPHSGD just adjusts the clipping threshold of parameters essentially, thus the difference between their expected errors is only a finite constant $C$. However, the expected error of LLAGP cannot be compared with other methods since the noise is injected into hidden layers instead of the tensors directly.
A comparison of the results in Table 5 shows that whether Laplacian noise or Gaussian noise is injected, the expected error of TLDP is $O(1/I)$, which is consistent with the conclusions of Theorem 2.
@@ 5.3.2 Type II Results
Fig. 2 presents the F1-score across different values of $\epsilon$, where the relaxation term $\delta$ is set to $10^{-5}$. Compared to Type I, the rate of increase in F1-score across the three experimental results slows down as $\epsilon$ increases in Type II. For the MNIST dataset, the F1-score reaches a maximum of about 80%, while in CIFAR-10 and SVHN, the increase is much more limited, with F1-scores hovering around 40% finally. Unlike in Type I, the performance of the Laplace, Gaussian, MVG, and IDN mechanisms in Type II shows occasional fluctuations in F1-scores, although they still remain at extremely low levels, indicating their ineffectiveness for classification tasks. Additionally, consistent with the theoretical error analysis presented earlier, the TLDP-G mechanism generally outperforms the TLDP-L mechanism in these experiments.
5.3.3 Type III Results
@@
Fig. 3 shows the results of Type III on the different datasets for comparison. Two new datasets, IMDB for text mining and DarkNet for data mining, and two new comparative methods, LLAGP and DPHSGD, are introduced in the experiment. In Type III, due to the transmission of more abstract model parameters to the central server, the results from the various methods only become comparable when $\epsilon$ exceeds 100. As anticipated, the performance of the Laplace, Gaussian, MVG, and IDN mechanisms remains significantly low. LLAGP outperforms all other comparative methods across most scenarios except on the CIFAR-10 dataset, but still fails to surpass the method proposed in this paper, while DPHSGD only delivers good performance on the Darknet dataset and performs poorly in all other cases. Overall, the TLDP method proposed in this paper achieves the best comprehensive performance.
@@
Fig. 3 displays the F-1 score for varying $\epsilon$ values, with the relaxation term $\delta$ fixed as $10^{-5}$. In Type III, the results for the MNIST, CIFAR-10, and SVHN datasets are consistent with those in Type I and Type II, with TLDP significantly outperforming the Laplace, Gaussian, MVG, IDN, and DPHSGD mechanisms. LLAGP performs best on the SVHN dataset, and can achieve accuracy close to that of TLDP when $\epsilon$ is sufficiently large; however, it performs worst on the CIFAR-10 dataset, with results that are not significantly different from those of other comparative experiments. On the IMDB dataset, the accuracy of the Laplace, Gaussian, MVG, and IDN mechanisms improves to around 40%, but no further improvement is observed as $\epsilon$ increases. In contrast, both TLDP methods consistently achieve accuracy levels of around 80%. LLAGP also performs well on this dataset, keeping pace with TLDP methods. On the DarkNet dataset, the performance of the different methods is more distinctly differentiated. The IDN and Laplace mechanisms show the poorest performance, followed by the Gaussian and MVG mechanisms. Although the TLDP method continues to demonstrate an advantage, the gap with other methods is less pronounced compared to the former experiments—especially since the accuracy of LLAGP and DPHSGD has ultimately become sufficiently close to that of TLDP. This is especially true for the TLDP-G mechanism, which initially exhibits the lowest accuracy when $\epsilon$ is set to 10 but eventually surpasses all other methods, aligning with the TLDP-L curve.
We find that the performance of IDN is not as good as claimed in [25], primarily due to its high reliance on its pre-defined utility subspace. When there is a lack of known conditions, its performance is significantly compromised. Most comparative methods (i.e., Gaplace, Gaussian, MVG, IDN) perform poorly in experiments, mostly because the amount of noise they inject is $O(I)$ times that of the TLDP method (where $I=784$ for the MNIST dataset and $I=1024$ for CIFAR-10 and SVHN datasets), and the larger value of $I$ results in the poorer performance of the comparative methods. It is worth noting that LLAGP performs better than Laplace, Gaussian, MVG, IDN, and DPHSGD; the underlying reason behind this could be that it selectively allocates less privacy budget to the hidden layers close to the output layer and more privacy budget to those that are close to the input layer to avoid privacy leakage (under its assumption that these hidden layers face a higher privacy leakage risk compared with the rest of the layers) instead of treat all hidden layers equally.
@@
The results from all the aforementioned experiments provide strong evidence of the robustness of the TLDP mechanism across diverse experimental settings. Regardless of the dataset or the tasks, TLDP consistently outperforms other baseline methods, showcasing its ability to adapt effectively to different environments. This robustness highlights TLDP’s potential as a reliable solution for preserving privacy in distributed computing systems while maintaining high model performance, making it a promising candidate for a wide range of distributed applications in real-world scenarios.
6 Conclusion
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (62406215) and the Science and Technology Program of Jiangsu Province (BZ2024062).